1 One box contains seven red balls and three green balls, and a second box contains six red balls and four green balls. A ball is randomly selected chosen from the first box and placed in the second box. Then a ball is randomly selected from the second box and placed in the first box. a. What is the probability that a red ball is selected from the first box and a red ball is selected from the second box? b. A the conclusion of the selection process, what is the probability that the numbers of red and green balls in the first box are identical to the numbers at the beginning
Answer #1

a)P(red ball from first box and red from second box)=P(red from first box)*P(red from second box|red from first box)

=(7/10)*(7/11)=49/110

b)

P(first box have identical numbers at the beginning)

=P(red from first box)*P(red from second box|red from first box)+P(green from first box)*P(green from second box|green from first box)

=(7/10)*(7/11)+(3/10)*(5/11)=64/110=32/55